20 October UT

18:20 - IIF problems - LBCstart had to be done since LBC code had been recompiled 18:30 - problems with sending a RB ob when red is unticked

Z11 = 1000 was entered by operator

19:10 - 021155 need manual correction, +1000 Z8 and -2000 Z7.

19:19 - 021700 still has too much coma for fpia, manual add -2000 Z7. 022046 -7656 ... z4 only 022348 2206 413 -712 -430 247 500 0.86 022620 223 -302 -197 -133 104 500 1.40 still adding a lot of Z11... 022808 48 200 -107 -486 193 262 1.38 023722 -118 17 -227 -1003 72 81 1.26 023955 896 -112 -1035 -2570 -355 -112 1.11

19:46 hitting negative Y limit but at only 63 deg elevation! So John & Aaron are using the WF GUI to adjust. T = 7.7deg C.

put in +20" tip which moved +0.9mm Y... WF gui zeroed AO offsets.

025029 - not fit by fpia - we need to add -2000 Z7 again. 025932 - another manual -2000 Z7. 030258 - -6486 030611 -7 -459 -887 259 991 93 0.86 030830 -14 99 -260 29 64 232 1.15 031112 -92 -228 232 58 -9 16 1.08

superfoc sequences pa=0: 031828 pa=180: ...032513 is last set.

For test of guiding, ngc7469guidetest.ob takes U-Bessel 180s x 5 dithers. L2 Z11

moved L2 by 1mm fastextra - needs lots of collim but eventually, best Z11 was around 140, in noise. Moved L2 by another 1mm

took superfoc sequence 042936...043236 followed by a fastextra (043515). moved L2 back to starting position 044104 - 2973 173 169 38 94 144 0.95 044313 - 49 -322 47 -70 0 -214 0.86

tweak lbcfpia to make an offset to the z11 - so the inner hole will be smaller on 'converged' images. 044313 - 49 -322 47 -70 0 -514 0.86

045848 - 27 103 -23 176 240 -363 0.75 050109 13 -157 -160 -240 -260 -172 0.95 050329 -74 135 -23 76 36 -93 1.01 050536 53 -138 -34 -30 -32 56 0.92 extra/intra focal pair with 4 chips: 050809 050905

John is updating the collimation table so that the start of night will not need manual adjustments.

recollim with fpia: 052435 -2228 -418 -796 100 51 -203 0.81 can we see any curvature in the images taken with L2 moved cf L2 not moved. pointing star ACT0375 should fall at rot center for blue: 1043, 2950 055835 actually at 1184, 2709 tried using ptadjust, but it seems to hang on 'click on pointing star'

going along with collimation now... image, ra, dec, elev, temp 01, 10, 67.24, 6.5 ...note only images which are collimated. 061759, 67.3358, 6.4C -54 93 -201 -15 -61 16 0.95

063218, 57.2750, 6.3C
  1. -203 293 -8 42 36 0.98 (astig did not go down well)

47deg elevation, hit -Y limit

* Global Offsets * X=1.463, Y=-1.72, Z=0.226, RX=-35.3, RY=-8.2, RZ=0

going back to 57 - but we are still at the limit.

going back to 67 - but now it is still worse... the coma makes y ao even more negative.

removed ao corrections...

070719, 64.9, 6.2,
  1. -351 -218 -272 80 23 1.03
but also what is 070901??

071718, 54.89, 6.0 21 -71 253 -45 30 -40 0.85 072305, 71.466, 5.9 -84 66 98 -194 93 -30 0.96

072909, 75.4, 6.1 -158 45 173 -68 30 69 0.84

073724, 84.12, 5.9
  1. -40 -101 -75 -110 10 0.82

changed Y, RX global offsets to 0.0, 0.0 to balance range of Y. this change was from Y=1.720 RX=-35.3

86.8 deg elevation 080045 --- bad focus --- did global z offset change???

go to lower elevation to check... 66deg - again focus is off...

completely lost with focus/collimation now ... note that the values John entered for start of night fix were x=0.7 and z=0.47. just in case we want to zero these.

083257 - lbcfpia - sent 083520 - 083801 - 98 -187 45 -367 -10 -87 0.72 at 65deg elevation, t=5.9C

084505 - -225 -227 -115 -19 -82 157 0.76 at 77.5 deg elevation, t=5.3C

85.66 deg - hit Y = 2.6 limit. 090348 -

56.9deg - -114 -401 -259 -84 27 -31 0.87 091330, t=5.8

39.3deg, -129 354 86 31 -168 54 1.08 092637, t=5.6

hit an X limit (X=2.6) for elevation 35deg.

62deg, -208 -305 26 -257 42 17 0.85 093653, t=5.6

86deg, very very close to Y=2.6

So we should stay between 35-85 degrees...

Get data for Andrew now:

collimate: 095824 -23 -279 -30 -156 81 -5 0.81 take extra/intra focal pair: 100022/100118 Z11 = 0 (as a reference...collimated)

Z11 = +500nm: 100341/100437 Z11 = +1000nm: 100648/100745 (another +500 but I clicked send 2x, adjust just 1)Z11 = -500nm: 101040/101136 (gave z11=-1500) Z11 = -1000nm: 101547/101643 (gave Z11=-500)

gave Z11=+1000nm to get back to nominal position and run fastextra again 102347 - 102534 171 -93 -234 74 46 0 0.79 102709 -15 -554 -145 -131 -99 140 1.05

superfoc pa=0 (again, since the earlier set for tilt was off in best-foc) 103004-103305 best-foc is 0.037, 2.78 pa=180... 103557-103857 best-foc is 0.037, 2.41

111150 111437 - 112055 -183 61 -199 -131 -7 -2 0.74 (still not converging too well, but go ahead and try) extra/intra focal pair (window of chip2) 112310, 112357

send +500 Z11 and take another extra/intra: 112711/112758 send another +500 - so now Z11= +1000 over nominal: 113019/113106 send Z11=-1500 to make Z11=-500 over nominal: 113344/113431 and another -500 to make Z11=-1000 over nominal: 113637/113724

add Z11=+1000 to return to nominal --- extra/intra - 114007/114054

-- OlgaKuhn - 20 Oct 2008
Topic revision: r1 - 21 Oct 2008, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback